x(squared)+3x-40=0
Answer by stanbon(75887)

x(squared)+3x-40=0
x(squared)+3x-40=0
x^2+3x-40=0
(x+8)(x-5)=0
x=-8 or x=5
-----------------
Cheers,
Stan H.